# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ VII Summary
We have presented a comprehensive survey of the theoretical and experimental status of about 200 rare and exclusive few-body decays of the four heaviest Standard Model (SM) particles: the Higgs, the electroweak (W, Z) bosons, and the top quark. Rare decays are defined here as those having branching fractions below $\mathcal{B}\approx 10^{-5}$, and we focus on those with two- to four-particles in the final state. Such decay processes remain experimentally unobserved and only upper limits have been set for about 50 of them. The study of these decay processes provides a powerful window into physics beyond the Standard Model (BSM) either directly, by probing SM suppressed or forbidden processes (such as flavour-changing neutral currents FCNC, lepton-flavour or lepton-flavour-universality violating processes, or spin-selection-rules violating decays), or indirectly as SM backgrounds to multiple exotic BSM decays (e.g., into axion-like particles ALPs, massive gravitons, or dark photons). Additionally, such decays offer a unique opportunity to constrain the light-quark Yukawa couplings, can provide alternative means to measure the W and/or top-quark masses, and help improve our understanding of quantum chromodynamics (QCD) factorization with small nonperturbative corrections.
@@
$\rm\gamma$ $+$ $\rm\rho^{0}$  $(1.68\pm 0.08)$$\times~{}10^{-5}$ $<$ $3.7$×$\times~{}10^{-4}$ [92] $\lesssim 5.7\times~{}10^{-5}$ $\sim$ 1/4
$\rm J/\psi$  $(2.95\pm 0.17)$$\times~{}10^{-6}$ $<$ $2.0$×$\times~{}10^{-4}$ [99] $\lesssim 3.9\times~{}10^{-5}$ [60] $\sim$ 1/10
$\rm W^{\mp}$ $+$ $\rm\rho^{\pm}$  $(1.5\pm 0.1)$$\times~{}10^{-5}$
@@
$\rm\omega$  $(2.82\pm 0.41)$$\times~{}10^{-8}$ $<$ $3.9$×$\times~{}10^{-6}$ [95] $\lesssim 5.8\times~{}10^{-7}$ $\sim$ 1/20
$\rm\phi$  $(1.17\pm 0.08)$$\times~{}10^{-8}$ $<$ $7.0$×$\times~{}10^{-7}$ [92] $\lesssim 5.0\times~{}10^{-8}$ $\sim$ 1/6
$\rm J/\psi$  $(9.96\pm 1.86)$$\times~{}10^{-8}$ $<$ $6.0$×$\times~{}10^{-7}$ [96] $\lesssim 3.1\times~{}10^{-7}$ [60] $\sim$ 1/3
$\rm\Upsilon(1S)$  $(4.93\pm 0.51)$$\times~{}10^{-8}$ $<$ $1.1$×$\times~{}10^{-6}$ [93] $\lesssim 1.7\times~{}10^{-7}$ $\sim$ 1/4
@@
$\rm D_{s}^{\pm}$  $\left(3.66^{+1.49}_{-0.85}\right)$$\times~{}10^{-8}$ $<$ $6.5$×$\times~{}10^{-4}$ [144] $\lesssim 1.2\times~{}10^{-5}$ $\sim$ 1/300
@@
W $+$ Z $+$   b $1.8$$\times~{}10^{-6}$
Future Circular Collider (FCC-ee and FCC-hh). From the number of H, Z, W bosons, and top quarks expected to be produced at the HL-LHC, FCC-ee, and FCC-hh colliders, and by statistically extrapolating the current 95% confidence-level limits set, we provide estimates of the conservatively achievable experimental upper bounds (or observations) in those machines. Among those, in Table 27 we have selected a few interesting channels that can be observed and/or deserve further study in p-p collisions at the HL-LHC. The last column indicates the approximate ratio of theoretical to expected experimental rates, $\mathcal{B}(\rm th)/\mathcal{B}(\rm exp)$. Two reasons motivate this selection of channels: either (i) they have relatively large rates, $\mathcal{O}(10^{-5})$ (with a measurement having been attempted or not, to date), or (ii) they have lower rates, $\mathcal{B}\lesssim 10^{-8}$, but measurements of upper bounds have been performed, and our conservative projections indicate $\mathcal{B}(\rm th)/\mathcal{B}(\rm exp)$ not far from unity. For the Higgs boson, we list first the rare $\rm H\to 4\gamma$ decay that has not been directly searched for at the LHC to date, but limits exist on the process $\rm H\to a(\gamma\gamma)a(\gamma\gamma)$ with two intermediate ALPs decaying into photons  [66, 88, 89, 90]. Although its rate makes its observation impossible at the LHC in the absence of enhancing BSM effects, it would not be too difficult for ATLAS/CMS to recast any current and future similar ALP searches into upper bounds on the $\rm H\to 4\gamma$ “continuum” decay. Next, we find that the HL-LHC can conservatively set upper bounds on the exclusive $\rm H\to\gamma\,\rho$ and $\rm H\to\gamma\,\mathrm{J}/\psi$ decays at about four and ten times their expected SM values, respectively. Similarly, we find that HL-LHC can potentially observe $\rm Z\to\gamma\,\omega,\,Z\to\gamma\,\mathrm{J}/\psi$, and $Z\to\gamma\,\Upsilon(1S)$ with ratios between the theoretical rates and projected experimental limits as large as 1/2, 1/3, and 1/4, respectively. In the W boson case, the least suppressed rare decays are $\rm W\to\gamma\,\pi,\,\gamma\,D_{\mathrm{s}}$, but with experimental bounds expected at about two orders-of-magnitude their SM values. Finally, the semiexclusive decay of the top quark into a charm jet and B meson (identified experimentally as a b-jet) has $\mathcal{O}(10^{-4})$ rates that make it observable at the HL-LHC [26], and the three-body $\rm t\to W\,Z\,b$ decay can occur in about one out of two million top-quark decays, and a search should also be attempted. Of course, the HL-LHC channel selection of Table 27 is driven by the SM rates and their potential visibility, but as explained in this work there are many other suppressed decays of the four heaviest particles that can be enhanced in many BSM scenarios, and should be an active part of target searches in the next years.

### v4

@@ VII Summary
We have presented a comprehensive survey of the theoretical and experimental status of about 200 rare and exclusive few-body decays of the four heaviest Standard Model (SM) particles: the Higgs, the electroweak (W, Z) bosons, and the top quark. Rare decays are defined here as those having branching fractions below $\mathcal{B}\approx 10^{-5}$, and we focus on those with two- to four-particles in the final state. Such decay processes remain experimentally unobserved and only upper limits have been set for about 50 of them. The study of these decay processes provides a powerful window into physics beyond the Standard Model (BSM) either directly, by probing SM suppressed or forbidden processes (such as flavour-changing neutral currents FCNC, lepton-flavour or lepton-flavour-universality violating processes, or spin-selection-rules violating decays), or indirectly as SM backgrounds to multiple exotic BSM decays (e.g., into axion-like particles ALPs, massive gravitons, or dark photons). Additionally, such decays offer a unique opportunity to constrain the light-fermion Yukawa couplings, can provide alternative means to measure the W and/or top-quark masses, and help improve our understanding of quantum chromodynamics (QCD) factorization with small nonperturbative corrections.
@@
[…]
$\rm W^{\mp}$ ++ $\rm\rho^{\pm}$ $(1.5\pm 0.1)$ $\times~10^{-5}$
@@
$\rm\phi$ $(1.17\pm 0.08)$ $\times~10^{-8}$ << $07\times~10^{-7}$ [ATLAS:2017gko] $\lesssim 5.0\times~10^{-8}$ $\sim$ 1/5
@@
[…]
W ++ Z ++   b $1.8$ $\times~10^{-6}$
Future Circular Collider (FCC-ee and FCC-hh). From the number of H, Z, W bosons, and top quarks expected to be produced at the HL-LHC, FCC-ee, and FCC-hh colliders, and by statistically extrapolating the current 95% confidence-level limits set, we provide estimates of the conservatively achievable experimental upper bounds (or observations) in those machines. Among those, in Table 27 we have selected a few interesting channels that can be observed and/or deserve further study in p-p collisions at the HL-LHC. The last column indicates the approximate ratio of theoretical to expected experimental rates, $\mathcal{B}(\rm th)/\mathcal{B}(\rm exp)$. Two reasons motivate this selection of channels: either (i) they have relatively large rates, $\mathcal{O}(10^{-5})$ (with a measurement having been attempted or not, to date), or (ii) they have lower rates, $\mathcal{B}\lesssim 10^{-8}$, but measurements of upper bounds have been performed, and our conservative projections indicate $\mathcal{B}(\rm th)/\mathcal{B}(\rm exp)$ not far from unity. For the Higgs boson, we list first the rare $\rm H\to 4\gamma$ decay that has not been directly searched for at the LHC to date, but limits exist on the process $\rm H\to a(\gamma\gamma)a(\gamma\gamma)$ with two intermediate ALPs decaying into photons  [ATLAS:2015rsn, CMS:2022fyt, CMS:2022xxa, ATLAS:2023ian]. Although its rate makes its observation impossible at the LHC in the absence of enhancing BSM effects, it would not be too difficult for ATLAS/CMS to recast any current and future similar ALP searches into upper bounds on the $\rm H\to 4\gamma$ “continuum” decay. Next, we find that the HL-LHC can conservatively set upper bounds on the exclusive $\rm H\to\gamma\,\rho$ and $\rm H\to\gamma\,\mathrm{J}/\psi$ decays at about four and ten times their expected SM values, respectively. Similarly, we find that HL-LHC can potentially observe $\rm Z\to\gamma\,\phi$, $\rm Z\to\gamma\,\mathrm{J}/\psi$, and $\rm Z\to\gamma\,\Upsilon(1S)$ with ratios between the theoretical rates and projected experimental limits as large as 1/5, 1/3, and 1/4, respectively. In the W boson case, the least suppressed rare decays are $\rm W\to\gamma\,\pi,\,\gamma\,D_{\mathrm{s}}$, but with experimental bounds expected at about two orders-of-magnitude their SM values. Finally, the semiexclusive decay of the top quark into a charm jet and B meson (identified experimentally as a b-jet) has $\mathcal{O}(10^{-4})$ rates that make it observable at the HL-LHC [dEnterria:2020ygk], and the three-body $\rm t\to W\,Z\,b$ decay can occur in about one out of two million top-quark decays, and a search should also be attempted. Of course, the HL-LHC channel selection of Table 27 is driven by the SM rates and their potential visibility, but as explained in this work there are many other suppressed decays of the four heaviest particles that can be enhanced in many BSM scenarios, and should be an active part of target searches in the next years.